OpenStudy (anonymous):

Solve the proportion for t. t/4.5 equals 14/9

OpenStudy (texaschic101):

t/4.5 = 14/9 cross multiply 9t = (4.5)(14) 9t = 63 t = 63/9 which reduces to 7 check.. 7/4.5 = 14/9 cross multiply (7)(9) = (4.5)(14) 63 = 63 (correct) t = 7

OpenStudy (anonymous):

t/4.5 = 14/9 t = (14/9) x 4.5 t = 7

OpenStudy (anonymous):

Thanks:)

OpenStudy (texaschic101):

:)

OpenStudy (anonymous):

welcome